# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 5.2.2 Identification with staggered designs
where $\overline{Y}_{i,t\leq g-1}=\sum_{s=1}^{g-1}Y_{i,s}\big{/}(g-1)$ is the time average of group $g$ pre-treatment periods for each unit $i$; see, e.g., Callaway2023_review, Lee_Wooldridge_2023, and the discussion in deChaisemartin2023-survey. Although the estimand does look different than the previous one, it is easy to see that it also resembles a $2\times 2$ design: it effectively leverages two periods, $t$ (post) and the average of all period $t<g$ (pre), and two groups, $G_{i}=g$ (treated) and $G_{i}>t$ (comparison). In practice, however, there is no general econometrics guarantee that estimators based on (5.8) will be more precise than estimators based on (5.6); see, e.g., Harmon2024. This arises because it is not always optimal to weigh all pre-treatment periods equally when forming estimators for $ATT(g,t)$. In fact, as discussed in Chen_SantAnna_Xie_effientDiD_2024, the (asymptotically) optimal way to aggregate information across pre-treatment periods and comparison groups under Assumption PT-GT-all depend on the correlation structure of how the outcome changes over time across different comparison groups. And an effective way to leverage this information consists of constructing DiD estimators for $ATT(g,t)$’s that efficiently weigh several $2\times 2$ DiD estimators for the $ATT(g,t)$ that use different comparison groups and different baseline periods. In sum, even these more complex DiD estimators closely resemble what we have done in the $2\times 2$ design.
@@
In the end, a natural question arises: Which kind of parallel trends assumption should one use? This context-specific question is hard to answer, as each assumption has pros and cons. For instance, Assumption PT-GT-all leads to more precise $ATT(g,t)$ estimators because it uses data from multiple pre-treatment periods and multiple comparison groups. As power is something important when conducting causal inference, this is appealing. On the other hand, it imposes parallel pre-trends, an assumption that is not required for identification of $ATT(g,t)$ and that we have not imposed in $2\times T$ DiD designs (see our discussion of equation (5.3)). If pre-trends are not parallel, then estimates of $ATT(g,t)$ based on Assumption PT-GT-all can be biased.
@@
The other extreme is to make Assumption PT-GT-Nev and only use comparison groups made up of never-treated units. This avoids compositional changes in the comparison group over time, does not restrict pre-trends, and identifies all the $ATT(g,t)$’s. [^28] Without never-treated units, we cannot estimate $ATT(g,t)$ for the last observed treatment date, which shapes the feasible target parameters. For example, in our Medicaid expansion example, without the presence of never-treated (by 2019) counties, we would not be able to estimate the treatment effects in 2019 (i.e., $ATT(g=2015,t=2019)$. On the other hand, never-treated units may have remained untreated for reasons related to $\Delta Y_{i,t}(0)$. Non-expansion counties may be too different from expansion counties for them to reflect the relevant counterfactual. This could be, in part, justified after examining the differences in covariate levels and trends between treatment and control groups, as we saw in the discussion of Table 4. Also, depending on how widespread treatment is, there may be too few never-treated units to obtain precise estimates.
@@
We view Assumption PT-GT-NYT as a middle step that uses all not-yet-treated units as a comparison group without restricting all pre-treatment trends to be parallel. It uses more information than Assumption PT-GT-Nev, which can lead to gains in precision and helps to incorporate covariates. While it uses less information than Assumption PT-GT-all, it is also less susceptible to bias from violations of parallel pre-trends. For our Medicaid application, we favor Assumption PT-GT-NYT, as we prefer to not impose parallel pre-trends from 2009 to 2014 or to rely exclusively on comparisons to the set of states that have not expanded Medicaid as of 2024.
@@ 5.2.3 Estimators for staggered designs without covariates
This simple estimator is what Callaway2021 proposes when one uses the not-yet-treated group as the comparison group.292929 Recently, Dube_etal_LPDID_2024 show that one can also get $ATT(g,t)$ estimates that are equivalent to $\widehat{ATT}_{\text{nyt}}(g,t)$ by using local projections. One can also get similar estimators using a “stacked DiD” procedure akin to what fadlon2021, deshpande2019, and Cengiz2019 have implemented.
@@
When Assumption PT-GT-all holds instead, one can construct plug-in estimators for (5.8), i.e.,
@@
Alternatively, Wooldridge2021 proposed to construct estimators for $ATT(g,t)$ based on Assumption PT-GT-all using following “extended” TWFE specification
@@
where $\beta^{W}_{g,t}$ are estimated using (weighted) least squares. Wooldridge2021 show that $\widehat{\beta}^{W}_{g,t}$ consistently estimate $ATT(g,t)$ under Assumption PT-GT-all, though we do not know the exact way that $\widehat{\beta}^{W}_{g,t}$ combines pre-treatment periods and not-yet-treated units, i.e., we do not know the statistical estimand associated with ${\beta}^{W}_{g,t}$. Wooldridge2021 also shows that $\widehat{\beta}^{W}_{g,t}$ is numerically the same as the “imputation” estimators proposed by Gardner2021, Liu2022, and Borusyak2023 with balanced panel data (and these specifications do not have covariates). [^30] Imputation procedures work in two-steps. The first step uses all untreated observations to run the regressions $Y_{it}=\theta_{t}+\eta_{i}+\epsilon_{it}$ using only data from the $(i,t)$ pairs that satisfy this criterion, and get the fitted values $\widehat{Y}_{i,t}(0)=\widehat{\theta}_{t}+\widehat{\eta}_{i}$ for all eventually-treated observations. The second step estimates $ATT(g,t)$ by $\widehat{ATT}_{\text{imp}}(g,t)=\frac{\sum_{i=1}^{n}\mathbf{1}\{G_{i}=g\}(Y_{i,t}-\widehat{Y}_{i,t}(0))}{\sum_{i=1}^{n}\mathbf{1}\{G_{i}=g\}}$. See Gardner2021, Liu2022, and Borusyak2023 for details. Note that the specification in (5.12) is similar to the Sun2021’s specification (5.11), but it omits the pre-treatment event-time dummies. This is justified because (5.12) imposes parallel pre-trends (Assumption PT-GT-all) while (5.11) does not (it effectively relies on Assumption PT-GT-Nev). Thus, in general, one should not expect $\widehat{\beta}_{g,t}^{W}$ to be equal to $\widehat{\beta}^{SA}_{g,t-g}$.
@@
Overall, these different $ATT(g,t)$ estimators highlight that we can leverage our DiD expertise built in the $2\times 2$ setup to estimate heterogeneity-rich parameters $ATT(g,t)$’s. The exact estimator we can use is shaped by the form of parallel trends. In our Medicaid application, we report in Figure 6 our $ATT(g,t)$ estimates based on Assumption PT-GT-NYT and (5.9). As we have four eventually-expanded sets of counties based on whether Medicaid expanded in 2014, 2015, 2016, or 2019, we report four sets of event studies, one for each expansion group. For the 2014, 2016, and 2019 expansion groups, we find that Medicaid did not lead to significant changes in adult mortality rates. For the 2015 expansion group, adult mortality rates rose after expansion. Regarding pre-trends, Figure 6 suggests that there may be some non-negligible pre-trends for the 2016 expansion group, though these are not statistically different from zero.
@@
In Section 4, we highlighted that unconditional parallel trends such as Assumption PT-GT-NYT may not be very plausible in our Medicaid context. A reason for this potential threat to our identification strategy is that important determinants of changes in untreated adult mortality rates are imbalanced across treatment and comparison groups. In such cases, one should interpret the results in Figure 6 with great care. In addition, it is also important to highlight that, as indicated in Table 1, the 2015, 2016, and 2019 expansion groups are relatively small and represent only 6%, 2%, and 3% of the US population, respectively. Thus, analyzing these groups separately may be “too noisy” and not representative of the overall effect for the US population. This does not mean that these $ATT(g,t)$’s are not useful. They are actually an essential part of our DiD analysis, but we may want to aggregate the cohorts to get more informative target parameters for the overall treated population. We now turn to how to aggregate the $ATT(g,t)$’s and then discuss how to incorporate covariates.
@@
Notes: This figure shows the group-time ATT estimates ($ATT(g,t)$) in calendar time for the four treatment timing groups of counties that expanded Medicaid before 2019, using not-yet-treated units as the comparison group, and their uniform confidence intervals at the 95% significance level. The outcome variable is the crude mortality rate for adults ages 20-64, and standard errors are clustered at the county level. The vertical line represents the year before Medicaid expansion (i.e., $g$ 1) for the timing group.
@@ 5.2.4 Aggregating group-time average treatment effects
Aggregation in staggered designs involves a notion of time (either calendar time $t$ or event-time $e=t-g$), a length of time (how many periods to aggregate across), and group weights (so larger treatment groups can “matter more” than smaller ones). Given some set of weights, it is simple to average the $ATT(g,t)$ building blocks into many kinds of summary parameters,
@@
where $w_{\omega,g,t}$ are “generic” ($\omega$-weighted) group and time-specific (non-negative) weights that sum up to one. Specific choices for the $w_{\omega,g,t}$ weights map to different ways to aggregate and present interpretable causal effects in this kind of complex setting.
@@
A basic observation about weighting is that we can only give positive weights to the $ATT(g,t)$’s that we actually identified and estimated. Earlier treated groups have $ATT(g,t)$ estimates for later event-times by definition (and later treated groups have estimates for earlier pre-trends), so it will not be possible to include every group in an aggregated event-study parameter at every event-time. To see which $ATT(g,t)$’s will contribute to our event study, Figure 7 recenters our $ATT(g,t)$ estimates in event time instead of calendar time. That is, we plot $ATT(g,g+e)$ against $e$ for each expansion group.
@@
Notes: This figure shows the group-time ATT estimates ($ATT(g,t)$) in relative event time for the four treatment timing groups of counties that expanded Medicaid before 2019, using not-yet-treated units as the comparison group. The outcome variable is the crude mortality rate for adults ages 20-64.
Based on Figure 7, we will take “vertical” weighted averages of each available $ATT(g,g+e)$ for each event time $e$. For instance, to estimate an aggregate event study in event time 0 (a measure of instantaneous treatment effects), we would average estimates of $ATT(2014,2014)$, $ATT(2015,2015)$, $ATT(2016,2016)$ and $ATT(2019,2019)$. When we are interested in event time 1, we would now average $ATT(2014,2015)$, $ATT(2015,2016)$, $ATT(2016,2017)$. The same logic applies to other event times.
@@
$ATT_{\text{es}}(e)$ $=\mathop{}\!\mathbb{E}_{\omega}\bigg{[}ATT(G,G+e)\bigg{|}G+e\in[2,T],G\leq T\bigg{]}$
@@
Note that $ATT_{\text{es}}(e)$ gives the average treatment effect among the units that have been exposed to treatment for exactly $e$ periods, conditional on being observed having participated in the treatment for that number of periods (the condition that $G+e\in[2,T]$) and ever-participating in the treatment by period $T$ ($G\leq T$). One can also take a simple average of all available post-treatment event times, $ATT_{\text{es}}(e),e\geq 0$, and report an overall ATT measure. See Callaway2021 for a discussion of alternative aggregations based on calendar time and groups.
Estimating $ATT_{\text{es}}(e)$ is straightforward and, once again, relies on the plug-in principle: we need to replace $ATT(g,g+e)$ with its sample analogs (that we already computed in Figure 6), and use the relative adult population share of expansion group $g$ among eventually-treated units as estimates of the event study weights. Figure 8 reports our population-weighted estimates of the event-study aggregation from event times -5 to 5, i.e., from 5 years before the Medicaid expansion until 5 years after the Medicaid expansion. We also report pointwise and simultaneous confidence intervals at the 95% significance level. Overall, the results suggest that Medicaid expansion has no effect on adult mortality rates among counties that eventually experience a Medicaid expansion. The pre-trends are also fairly close to zero, suggesting that our parallel trends assumption may be reasonable.
@@
Notes: This figure shows the event study estimates with staggered treatment timing using the doubly-robust estimation method from Callaway2021 using not-yet-treated units as the comparison group. The outcome variable is the crude mortality rate for adults ages 20-64. The point estimate is reported by the circles, and both 95% point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines. We also report the simple average of all non-negative event times as a summary of the overall ATT (together with their standard errors and 95% confidence interval).
We conclude this section by stressing that the way we have constructed the event study parameters in Figure 8 uses all available information from Figure 6. A potential drawback of this strategy is that we do not always use the same set of groups across all event times. Practitioners usually refer to this as imbalance in event time. For instance, the 2019 expansion group only contributes to event time $e=0$, but not for $e=1$ or later event times. When compositional changes are a concern, one can impose balance in event time and estimate a balanced event study aggregation:
$ATT_{\text{es,bal},[\underline{e},\overline{e}]}(e)$ $=\mathop{}\!\mathbb{E}\bigg{[}ATT(G,G+e)\bigg{|}G+\overline{e}\in[2,T],G+\underline{e}\in[1,T],G\leq T\bigg{]}$
@@
$w_{g,[\underline{e},\overline{e}]}^{es,bal}=\mathbf{1}\{g+\overline{e}\leq T\}\mathbf{1}\{g+\underline{e}\geq 1\}P_{\omega}(G=g|G+\overline{e}\in[2,T],G+\underline{e}\in[1,T],G\leq T).$

### v2

@@ 5.2.2 Identification with staggered designs
where $\overline{Y}_{i,t\leq g-1}=\sum_{s=1}^{g-1}Y_{i,s}\big{/}(g-1)$ is the time average of group $g$ pre-treatment periods for each unit $i$; see, for example, Callaway2023_review, Lee_Wooldridge_2023, and the discussion in deChaisemartin2023-survey. Although the estimand does look different from the previous one, it too resembles a $2\times 2$ design: it effectively leverages two periods, $t$ (post) and the average of all period $t<g$ (pre), and two groups, $G_{i}=g$ (treated) and $G_{i}>t$ (comparison). In practice, however, there is no general econometrics guarantee that estimators based on (5.8) will be more precise than estimators based on (5.6); see, for instance, Harmon2024. This arises because it is not always optimal to weigh all pre-treatment periods equally when forming estimators for $ATT(g,t)$. In fact, as discussed in Chen_SantAnna_Xie_effientDiD_2024, the optimal way (or, more formally, the semiparametric efficient way) to aggregate information across pre-treatment periods and comparison groups under Assumption PT-GT-all depends on the correlation structure of how the outcome changes over time across different comparison groups. And an effective way to leverage this information consists of constructing DiD estimators for $ATT(g,t)$’s that efficiently weigh several $2\times 2$ DiD estimators for the $ATT(g,t)$ that use different comparison groups and different baseline periods. These efficiency weights do not depend on additional hard-to-motivate assumptions (e.g., homoskedasticity or restriction on the serial correlation), arise as a consequence of the information content of the identification assumptions, and can be transparently visualized using the tools provided Chen_SantAnna_Xie_effientDiD_2024. What is perhaps most important here is that even these more complex DiD estimators closely resemble the approach we took in the $2\times 2$ design.
@@
In the end, a natural question arises: Which parallel trends assumption should one use? This context-specific question is hard to answer, as each assumption has pros and cons. For instance, Assumption PT-GT-all leads to more precise $ATT(g,t)$ estimators because it uses data from multiple pre-treatment periods and multiple comparison groups. Given that power is important when conducting causal inference, this is appealing. On the other hand, it imposes parallel pre-trends, an assumption that is not required for identification of $ATT(g,t)$ and that we have not imposed in $2\times T$ DiD designs (see our discussion of equation (5.3)). If pre-trends are not parallel, then estimates of $ATT(g,t)$ based on Assumption PT-GT-all can be biased.
@@
The other extreme is to make Assumption PT-GT-Nev and use only comparison groups made up of never-treated units. This avoids compositional changes in the comparison group over time, does not restrict pre-trends, and identifies all the $ATT(g,t)$’s. [^27] Without never-treated units, we cannot estimate $ATT(g,t)$ for the last observed treatment date, which shapes the feasible target parameters. For example, in our Medicaid expansion example, without the presence of never-treated (by 2019) counties, we would not be able to estimate the treatment effects in 2019 (i.e., $ATT(g=2015,t=2019)$. It also avoids using as a comparison group units that may have chosen to begin treatment in a given period because of pre-treatment outcomes, which potentially violates parallel trends. For instance, states that expanded Medicaid in 2016 may have done so on the basis of county mortality rates from previous years, likely violating the parallel trends assumption for this group. On the other hand, never-treated units may have remained untreated for reasons related to trends in $Y_{i,t}(0)$. Non-expansion counties may be too different from expansion counties for them to reflect the relevant counterfactual. This could be, in part, justified after examining the differences in covariate levels and trends between treatment and control groups, as we saw in the discussion of Table 4. Also, depending on how widespread treatment is, there may be too few never-treated units to obtain precise estimates.
@@
We view Assumption PT-GT-NYT as a middle step that uses all not-yet-treated units as a comparison group without restricting all pre-treatment trends to be parallel. It uses more information than Assumption PT-GT-Nev, which can lead to gains in precision and helps to incorporate covariates. While it uses less information than Assumption PT-GT-all, it is also less susceptible to bias from violations of parallel pre-trends. For our Medicaid application, we favor Assumption PT-GT-NYT, as we prefer not to impose parallel pre-trends from 2009 to 2014 or to rely exclusively on comparisons to the set of states that have not expanded Medicaid as of 2024. On the other hand, if parallel trends are not plausible for a particular group of eventually-treated units, perhaps owing to selection based on time-varying unobservables (Ghanem2023), it is important to remove these units from the DiD analysis to retain interpretability. Ultimately, the plausibility of each parallel trends assumption may vary across different contexts. At the very least, we strongly recommend that researchers clearly state the specific parallel trends assumption they are actually imposing in their analysis to allow readers to discuss its plausibility in a scientifically grounded manner.
@@ 5.2.3 Estimators for staggered designs without covariates
This simple estimator is what Callaway2021 propose when one uses the not-yet-treated group as the comparison group.282828 Recently, Dube_etal_LPDID_2024 show that one can also get $ATT(g,t)$ estimates that are equivalent to $\widehat{ATT}_{\text{nyt}}(g,t)$ by using local projections. One can also get similar estimators using a “stacked DiD” procedure akin to what fadlon2021, deshpande2019, and Cengiz2019 have implemented.
@@
When Assumption PT-GT-all holds instead, one can construct plug-in estimators for (5.8):
@@
Alternatively, Wooldridge2021 proposed constructing estimators for $ATT(g,t)$ based on Assumption PT-GT-all, using following “extended” TWFE specification:
@@
where the $\beta^{W}_{g,t}$’s are estimated using (weighted) least squares. Wooldridge2021 shows that $\widehat{\beta}^{W}_{g,t}$ consistently estimate $ATT(g,t)$ under Assumption PT-GT-all, though we do not know the exact way that $\widehat{\beta}^{W}_{g,t}$ combines pre-treatment periods and not-yet-treated units, which is to say that we do not know the statistical estimand associated with ${\beta}^{W}_{g,t}$. Wooldridge2021 also shows that $\widehat{\beta}^{W}_{g,t}$ is numerically the same as the “imputation” estimators proposed by Gardner2021, Liu2022, and Borusyak2023 with balanced panel data (and these specifications do not have covariates). [^29] Imputation procedures work in two-steps. The first step uses all untreated observations to run the regressions $Y_{it}=\theta_{t}+\eta_{i}+\epsilon_{it}$ using only data from the $(i,t)$ pairs that satisfy this criterion, and get the fitted values $\widehat{Y}_{i,t}(0)=\widehat{\theta}_{t}+\widehat{\eta}_{i}$ for all eventually-treated observations. The second step estimates $ATT(g,t)$ by $\widehat{ATT}_{\text{imp}}(g,t)=\frac{\sum_{i=1}^{n}\mathbf{1}\{G_{i}=g\}(Y_{i,t}-\widehat{Y}_{i,t}(0))}{\sum_{i=1}^{n}\mathbf{1}\{G_{i}=g\}}$. See Gardner2021, Liu2022, and Borusyak2023 for details. Note that the specification in (5.12) is similar to the Sun2021’s specification (5.11), but it omits the pre-treatment event-time dummies. This is justified because (5.12) imposes parallel pre-trends (Assumption PT-GT-all) while (5.11) does not (it effectively relies on Assumption PT-GT-Nev). Thus, in general, one should not expect $\widehat{\beta}_{g,t}^{W}$ to be equal to $\widehat{\beta}^{SA}_{g,t-g}$.
@@
Overall, these different $ATT(g,t)$ estimators highlight that we can leverage our DiD expertise built in the $2\times 2$ setup to estimate heterogeneous $ATT(g,t)$ parameters. The exact estimator we can use is shaped by which parallel trends assumption holds. In our Medicaid application, we report in Figure 6 our $ATT(g,t)$ estimates based on Assumption PT-GT-NYT and (5.9). As we have four sets of counties defined by their Medicaid expansion timing, we report four sets of event studies, one for each expansion group. For the 2014, 2016, and 2019 expansion groups, we find that Medicaid did not lead to significant changes in adult mortality rates. For the 2015 expansion group, adult mortality rates rose after expansion. Regarding pre-trends, Figure 6 suggests that there may be some non-negligible pre-trends for the 2016 expansion group, though these are not statistically different from zero.
@@
In Section 4, we highlighted that unconditional assumptiosn such as Assumption PT-GT-NYT may fail in our Medicaid context. One reason is that important determinants of changes in untreated adult mortality rates are imbalanced across treatment and comparison groups. In such cases, one should interpret the results in Figure 6 with great care. In addition, it is also important to highlight that, as indicated in Table 1, the 2015, 2016, and 2019 expansion groups are relatively small and represent only 6%, 2%, and 3% of the US population, respectively. Thus, analyzing these groups separately may be “too noisy” and not representative of the overall effect for the US population. This does not mean that these $ATT(g,t)$’s are not useful. They are actually an essential part of our DiD analysis, but we may want to aggregate the cohorts to get more informative target parameters for the overall treated population. We now turn to how to aggregate the $ATT(g,t)$’s and then discuss how to incorporate covariates.
@@
Notes: This figure shows the group-time ATT estimates ($ATT(g,t)$) in calendar time for the four groups of counties that expanded Medicaid before 2019. Each panel uses 1,222 not-yet-treated counties as the comparison group and shows their uniform confidence intervals at the 95% significance level. There are 978 counties in states in the 2014 expansion group, 171 counties in the 2015 expansion group, 93 counties in the 2016 expansion group, 140 counties in the 2019 expansion group. The outcome variable is the crude mortality rate for adults ages 20-64, and standard errors are clustered at the county level. The vertical line represents the year before Medicaid expansion (i.e., $g$ 1) for the timing group. All results use population weights.
@@ 5.2.4 Aggregating group-time average treatment effects
Aggregation in staggered designs involves a notion of time (either calendar time $t$ or event-time $e=t-g$), a length of time (how many periods to aggregate across), and group weights (so larger treatment groups can “matter more” than smaller ones). Given some set of weights, it is simple to average the $ATT(g,t)$ building blocks into many kinds of summary parameters:
@@
where $w_{\omega,g,t}$ is a “generic” ($\omega$-weighted) group and time-specific (non-negative) weight that sums up to one. Specific choices for the $w_{\omega,g,t}$ weights map to different ways to aggregate and present interpretable causal effects in this kind of complex setting.
@@
A basic observation about weighting is that we can give positive weights only to the $ATT(g,t)$’s that we actually identified and estimated. Earlier treated groups have $ATT(g,t)$ estimates for later event-times by definition (and later treated groups have estimates for earlier pre-trends), so it will not be possible to include every group in an aggregated event-study parameter at every event-time. To see which $ATT(g,t)$’s will contribute to our event study, Figure 7 recenters our $ATT(g,t)$ estimates in event time instead of calendar time. That is, we plot $ATT(g,g+e)$ against $e$ for each expansion group.
@@
Notes: This figure shows the group-time ATT estimates ($ATT(g,t)$) in relative event time for the four treatment timing groups of counties that expanded Medicaid before 2019, using not-yet-treated units as the comparison group. The outcome variable is the crude mortality rate for adults ages 20-64. All estimates use population weights.
We will take “vertical” weighted averages of each available $ATT(g,g+e)$ for each event time $e$ based on Figure 7. For instance, to estimate an aggregate event study in event time 0 (a measure of instantaneous treatment effects), we would average estimates of $ATT(2014,2014)$, $ATT(2015,2015)$, $ATT(2016,2016)$ and $ATT(2019,2019)$. When we are interested in event time 1, we would now average $ATT(2014,2015)$, $ATT(2015,2016)$, $ATT(2016,2017)$. The same logic applies to other event times.
@@
$ATT_{\text{es}}(e)$ $=\mathop{}\!\mathbb{E}_{\omega}\bigg{[}ATT(G,G+e)\bigg{|}G+e\in[1,T],G\leq T\bigg{]}$
@@
Note that $ATT_{\text{es}}(e)$ gives the average treatment effect among the units that have been exposed to treatment for exactly $e$ periods, conditional on being observed having participated in the treatment for that number of periods (the condition that $G+e\in[1,T]$) and ever-participating in the treatment by period $T$ ($G\leq T$). One can also take a simple average of all available post-treatment event times, $ATT_{\text{es}}(e),e\geq 0$, and report an overall ATT measure. See Callaway2021 for a discussion of alternative aggregations based on calendar time and groups.
Estimating $ATT_{\text{es}}(e)$ is straightforward and, once again, relies on the plug-in principle: we need to replace $ATT(g,g+e)$ with its sample analogs (which we already computed in Figure 6), and use the relative adult population share of expansion group $g$ among eventually-treated units as estimates of the event study weights. Figure 8 reports our population-weighted estimates of the event-study aggregation from event times -5 to 5, which respectively correspond to 5 years before the Medicaid expansion and 5 years after the Medicaid expansion. We also report pointwise and simultaneous confidence intervals at the 95% significance level. Overall, the results suggest that Medicaid expansion has no effect on adult mortality rates among counties that eventually experience a Medicaid expansion. The pre-trends are also fairly close to zero, suggesting that our parallel trends assumption may be reasonable.
@@
Notes: This figure shows the event study estimates with staggered treatment timing using the doubly-robust estimation method from Callaway2021, using 1,222 not-yet-treated units as the comparison group. The sample sizes used to estimate ATT parameters for each timing group are 2,200 for the 2014 group, 1,393 for the 2015 group, 1,315 for the 2016 group, and 1,362 for the 2019 group. The outcome variable is the crude mortality rate for adults ages 20-64. The point estimate is reported by the circles, and both 95% point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines. We also report the simple average of all non-negative event times as a summary of the overall ATT (together with their standard errors and 95% confidence interval). All results use population weights.
We conclude this section by stressing that the way we have constructed the event study parameters in Figure 8 uses all available information from Figure 6. A potential drawback of this strategy is that we do not always use the same set of groups across all event times. Practitioners usually refer to this as “imbalance in event time.” For instance, the 2019 expansion group contributes only to event time $e=0$, not $e=1$ or later event times. When compositional changes are a concern, one can impose balance in event time and estimate a balanced event study aggregation:
$ATT_{\text{es,bal},[\underline{e},\overline{e}]}(e)$ $=\mathop{}\!\mathbb{E}\bigg{[}ATT(G,G+e)\bigg{|}G+\overline{e}\in[1,T],G+\underline{e}\in[1,T],G\leq T\bigg{]}$
@@
$w_{g,[\underline{e},\overline{e}]}^{es,bal}=\mathbf{1}\{g+\overline{e}\leq T\}\mathbf{1}\{g+\underline{e}\geq 1\}P_{\omega}(G=g|G+\overline{e}\in[1,T],G+\underline{e}\in[1,T],G\leq T).$